OpenStudy (anonymous):

1) Solve the following system of linear equations using the substitution method. Be sure to explain all steps: -x + y = 3 2x + y = 9

OpenStudy (kamibug):

\(\huge\cal\color{royalblue}{Welcome~to~Openstudy!~:)}\) Step 1. Solve the first equation for y. y = 3 + x Step 2. Substitute that value in for y in the second equation and solve for x. 2x + (3 + x) = 9 3x + 3 = 9 3x = 6 x = 2 <--- Step 3. Substitute the value of x into the first equation and solve for y. -(2) + y = 3 y = ? <---

OpenStudy (anonymous):

y =-1.5

OpenStudy (kamibug):

Actually, you want to add 2 to both sides to finish solving for y. :) -2 + 2 + y = 3 + 2 0 + y = 3 + 2 y = 3 + 2 y = ?

OpenStudy (anonymous):

oh 5

OpenStudy (kamibug):

You've got it! ^_^

OpenStudy (anonymous):

awesome thank you so much!!!

OpenStudy (kamibug):

No problem :D